Consider the following gas phase reaction. D B—A=A—B + 2 D-D D CA + ť be B-B Consider the following bond energies. ASA: 600.Options for Part b : 1.) negative 2.)  postive or 3.) nearly zero

Options for Part C: 1.) Not very extensive at low temps but more so at high temps or 2.) extensive at all temps with only slight variations in temperature 3.) extensive at low temps but more so at high temps or 4.) extensive and its extent would not vary much with temperature or 5.)  extensive at low temps but less so at high temps or 6.) Not very extensive at low temps but even less so at high temps
Answer #1

tay the reaction is B-A=A-B + 2D-D - A=A & B-B the enthalpy of the reaction is given by A Hoxu - AH(B-B) + 4H (A=A) + 4 AH (D